OpenStudy (anonymous):

In the equation y = 2x2 - 7x - 9, which of the following is a value of x when y = 0? A. 2 1/4 B. -9 C. 4 1/2 D. 3

OpenStudy (anonymous):

find the roots of the equation when y=0

OpenStudy (anonymous):

2x2-7x-9=0 2x2+2x-9x-9=0 2x(x+1)-9(x+1)=0 (2x-9)(x+1)=0 x=9/2 or -1 hence answer C

OpenStudy (anonymous):

thx